Stat 50 Final ExamLast Name First Name 13. Consider the following paired data involving the systolic and diastolic blood pressure readin Systolic 125 120 130 130 144 143 140 130 Diastolic 82 91 100 100 80 90 80 80 8098 105 85 70 100 138 130 135 140 120 (a) Does there appear to be a correlation between the systolic and diastolic blood pressure? What is the regression line? Find the best predicted diastolic blood pressure for a person with a systolic reading of 122 (e) 14. Give two examples that we use ANOVA.
Answer #1

a)

Scatterplot of Diastolic vs Systolic 110 100 9 90 80 70 140 145 150 135 Systolic 120 125 130

From the above scatter plot, we can observe that increase the value of systolic increases the value of diastolic and vice versa. Hence, we may conclude that these two variable has a positive correlation.

b)

Regression Analysis: Diastolic versus Systolic

The regression equation is
Diastolic = 2.32 + 0.6374 Systolic


Predictor Coef SE Coef T P
Constant 2.32 36.06 0.06 0.950
Systolic 0.6374 0.2675 2.38 0.033


S = 9.02937 R-Sq = 30.4% R-Sq(adj) = 25.1%

The regression line is

Diastolic = 2.32 + 0.6374 Systolic

c)

The best predicted diastolic blood pressure for a person with a systolic reading of 122 is

Diastolic = 2.32 + 0.6374 *122 = 80.0828.

Q 4) Two example of the ANOVA table is:

1) Regression model: In a multiple regression model, the number of the independent variable is more than one. In this case, we use the ANOVA table to test the significance of at least one slope of the corresponding independent variable. If is significant we continue to test the significance of separate slopes.

2) To compare more than two means: For test more than two means, the z-test and t-test are not implacable. Hence, we use the ANOVA to test the significance of at least two means are significantly different. If it is significant, we proceed the pair-wise mean test like Bonferroni, Tukey etc.